# Lightning declines over shipping lanes following regulation of fuel sulfur emissions

### v3

@@ Introduction
By acting as cloud condensation nuclei (CCN), aerosol particles influence clouds and, in turn, the Earth’s energy balance. These aerosol interactions with clouds represent a significant uncertainty in our understanding of the Earth’s climate (Boucher et al.,, 2013). Shipping lanes can be used as a natural experiment to reduce that uncertainty. Fuel combustion by maritime shipping vessels in the open ocean leads to the emission of aerosol particles and associated precursors, such as SO2, into relatively clean marine air. These emissions can perturb low-level marine stratus cloud droplet number distributions and related macrophysical properties, such as cloud albedo and lifetime, by increasing CCN concentrations.(Diamond et al.,, 2020; Yuan et al.,, 2022; Durkee et al.,, 2000; Radke et al.,, 1989).
@@
Deep convective cloud (DCC) systems, commonly known as cumulonimbus or thunderstorms when lightning is generated, occur throughout the tropics, and are essential to the Earth’s water and energy cycles. In the tropics and some mid-latitude regions, the majority of precipitation and extreme weather is associated with DCCs (Feng et al.,, 2021). However, there is no strong consensus on the mechanisms or magnitudes of aerosol particle impacts on DCCs (Tao et al.,, 2012; Seinfeld et al.,, 2016; Igel and van den Heever,, 2021; Varble et al.,, 2023). Thornton et al (Thornton et al.,, 2017) documented a potential case of persistent maritime aerosol-DCC interactions analogous to the stratocumulus ship tracks, with the discovery of enhancements in lightning located over the two busiest shipping lanes in the world, which pass through the Indian Ocean and South China Sea (Fig. 1). Lightning results from cloud electrification, which in turn requires sufficient updraft velocity, vertical ice fluxes, and super-cooled liquid water in the mixed-phase region of a DCC to induce and sustain charge separation (Takahashi and Miyawaki,, 2002; Deierling et al.,, 2008). Several relevant mechanisms have been proposed by which aerosol particle could enhance lightning frequency over the polluted shipping corridors, all of which involve additional aerosol particles from ship emissions causing enhanced cloud droplet nucleation (Twomey,, 1977), which subsequently leads to either 1) a perturbation to the super-cooled liquid water and ice hydrometeor distributions in the mixed-phase region of DCC (Mansell and Ziegler,, 2013; Blossey et al.,, 2018; Sun et al.,, 2024); or 2) an increase in the frequency or intensity of deep convection due to changes in the vertical distribution of humidity (Abbott and Cronin,, 2021) or heating (Fan et al.,, 2018; Grabowski and Morrison,, 2020). Some combination of 1 or 2 is also possible.
@@
In January 2020, the International Maritime Organization (IMO) reduced the amount of allowable sulfur in fuel by a factor of seven, from 3.5% to 0.5% to curb effects of the maritime shipping industry on air pollution (IMO,, 2020). After similar regulations of fuel sulfur content, a study of ship plumes in the Baltic Sea documented a shift in the aerosol size distribution to smaller sizes (Seppälä et al.,, 2021). Recent analyses of shallow stratocumulus marine clouds over shipping lanes find changes to cloud brightness, droplet number, and droplet size, amounting to a globally averaged radiative forcing perturbation of O(0.1 Wm-2) associated with the IMO regulation, presumably due to the shift in aerosol size distribution (Watson-Parris et al.,, 2022; Yuan et al.,, 2022; Diamond,, 2023).
@@
Given that shallow clouds have exhibited a shift in droplet size distribution since the IMO regulation of sulfur, so too might DCCs, with downstream effects on mixed-phase microphysics, dynamics, and lightning. Here, we investigate how recent changes in ships’ fuel sulfur content impact lightning over the shipping lanes in the tropical Indian Ocean and South China Sea. We show that the shipping lane lightning enhancement decreases significantly with the onset of the IMO regulation and that this decrease in lightning is persistent in time across a range of atmospheric conditions. We further show that the mean cloud droplet number concentration of shallow warm clouds over the Indian Ocean shipping lane was enhanced before the IMO regulation and also exhibits a decrease since the IMO regulation. Finally, we discuss the implications of these results for the mechanisms of shipping lane lightning enhancement and for aerosol particle-DCC interactions generally.
@@
The Port of Singapore accounts for roughly 20% of the world’s bunkering fuel demand, and the two primary shipping lanes it services—the Indian Ocean and South China Sea (hereafter ”the shipping lanes”)—have nearly an order of magnitude higher traffic than other shipping lanes around the world (Figure 1, top panel) (Mao et al.,, 2022). We show that there is an enhancement in mean absolute lightning stroke density measured by the World Wide Lightning Detection Network (WWLLN) associated with these shipping lanes, consistent with previous work of Thornton et al, updated here for the time period of 2010 to 2019 in Figure 1 (middle panel). Since 2020, when the IMO regulation of sulfur emissions began, absolute lightning stroke density over the shipping lanes has declined by about 1 stroke km-2 year-1 as indicated by the difference in lightning stroke density between the mean of years 2020-2023 and that of 2010-2019 shown in Figure 1 (bottom panel). While some of the largest absolute declines in lightning since 2020 occur over the shipping lanes, lightning has increased or decreased in other parts of this region as well. As we illustrate below, changes in the dynamic and thermodynamic context for convection over these shipping lanes, as well as potential changes to shipping due to pandemic and other disruption to trade must be taken into account by controlling for variability in the amount and intensity of convection that leads to lightning as well as proxies for ship traffic.
Ship traffic in the vicinity of Singapore, as measured by vessel fuel sales at the Port of Singapore, has been relatively constant or even increased since 2020 (Figure S1) (Port of Singapore,, 2024). The disruption by COVID-19 did not obviously decrease activity at the port, seeming only to have briefly slowed the growth of cargo throughput for 2-3 months in 2020 (Gu et al.,, 2023). Therefore, we focus on controlling for the variability in background meteorological conditions that impact the frequency and intensity of convection, and thus lightning, over the shipping lanes.
@@
We first examine changes in the lightning stroke density enhancement over the shipping lanes using two basic controls on background meteorology: 1) we only sample from times and pixels with precipitating clouds, using IR and radar-calibrated microwave precipitation measurements from integrated multi-satellite retrievals (IMERG) (Huffman et al.,, 2015; Pradhan and Markonis,, 2023; Watters et al.,, 2023); and 2) we restrict analyses to the specific seasons in each region when conditions for lighting are favorable (November to April in the Indian Ocean; June to November in the South China Sea. Using these two criteria, we composite the lightning observations as a function of distance to the shipping lanes, the center of which we define as the peak in shipping emissions from the EDGAR emissions inventory (see Methods). As shown in Figure 2, mean absolute lightning as a function of distance from the shipping lanes exhibits a clear enhancement over the shipping lane before (pre-IMO), between approximately 150km south to 150km north of the shipping lanes, that has decreased since the regulation (post-IMO) (Figure 2a). Consistent with Figure 1, lightning has also fallen region-wide, particularly to the south of the shipping lanes. Therefore, we next examine whether the change in absolute lightning observed over the shipping lane since 2020 is greater than expected from natural sources of variability.
@@
To account for inter-annual variability in the frequency and intensity of convection in the region, we construct an estimate of the natural (unperturbed) annual lightning by regressing the observed annual lightning at a given distance from the shipping lane against three variables known to relate to lightning stroke density: Convective Available Potential Energy (CAPE, discussed further below) and precipitation rates (Romps et al.,, 2018), as well as the annual mean Oceanic Niño Index (ONI), subjected to the same two controls listed above — non-zero precipitation (except ONI) during the high lightning season. The influence of El Niño is assumed to be uniform across the domain for a given year. Inter-annual variability in the MJO was small and had a negligible impact when included in the regression (see SI). The regressed variables explain roughly 33% of variance of the annual means. The regressed predictions for lightning are then subtracted from the observed annual mean, leaving the anomalous mean lightning stroke density, i.e., that which cannot be explained by interannual variability in storm occurrence and intensity, the time series of which is shown as a function of distance from the shipping lane in Figure 2b. The reader is referred to Appendix A for further details about the regression.
@@
Having regressed out natural variations in meteorological drivers of lightning, the annual anomaly in lightning over the shipping lanes prior to 2020 is even more clear, as is the near step-change decrease in the anomaly after 2020. On the right axis of Figure 2b w also show the monthly fraction of fuel sales at the Port of Singapore that are high- and low-sulfur. Prior to the IMO regulation, essentially 100% of fuel sold at the port was the high sulfur type; correspondingly, the lightning anomaly over the shipping lane was 4.0 strokes km-2 year-1 on average and never fell below 3 strokes km-2 year-1 for more than one year at a time. Then, as evidenced by the bunkering fuel sales, adoption of the new regulation was immediate in 2020, such that high sulfur fuel has comprised less than 35% of fuel sold at the Port of Singapore. Total absolute fuel sales have increased since 2020, consistent with overall higher ship traffic (Figure S1); the Port of Singapore also experienced very little attenuation of sales associated with the onset of COVID-19, as noted by studies of port operations (Gu et al.,, 2023). Since the adoption of lower-sulfur fuels in January 2020, the shipping lane lightning enhancements have declined significantly: the anomaly has fallen to 1.8 strokes km-2 year-1 on average, with none of the recent 4 years reaching higher than 2.2 strokes km-2 year-1. That is, the enhancement in lightning over the shipping lane compared to adjacent regions has declined by approximately 50%.
@@
The preceding analysis relies on an assumption that the annual mean CAPE, precipitation, and ONI account for year-to-year variations in the regional drivers of convection and lightning. To further control for higher-frequency natural variations in convective activity and intensity, we examine the lightning enhancement in a 2-dimensional CAPE and precipitation space, using 3-hourly coincident observations of CAPE, precipitation, and lightning. Romps et al., (2018) showed that CAPE $\times$ precipitation is a reasonable proxy for lightning frequency over land, and Cheng et al., (2021) showed that, after adjusting for a CAPE threshold, CAPE $\times$ precipitation is a reasonable proxy for tropical oceanic lightning frequency. CAPE provides an estimate of the energetic potential for deep convection and the associated updraft strength while precipitation rate indicates both the presence of a storm and a measure of its intensity. By partitioning 3-hourly lightning observations into CAPE and precipitation bins (see Figure 3), we examine how regional and temporal variations in the environmental conditions impact the shipping lane lightning stroke density enhancement. Insofar as they are relevant to lightning generation, 3-hourly CAPE and precipitation observations capture variability from more indirect sources, such as sea surface temperatures (SST), MJO, fronts, etc (see SI for further discussion).
We compute lightning frequency in each CAPE-Precipitation bin using data from a region centered over each shipping lane and from reference regions adjacent to the shipping lanes, similar to those in (Thornton et al.,, 2017) (see Figure S2). We then compute a relative enhancement in lightning over the shipping lanes, before and after the onset of the IMO regulation, by taking the difference between corresponding CAPE-Precipitation bin-means in the shipping lane and associated reference box. The resulting shipping lane lightning enhancements as a function of both CAPE and Precipitation are shown in Figure 3. Before the IMO regulation (Pre-IMO), a strong shipping lane lightning enhancement existed in nearly every thermodynamic setting (e.g., in each CAPE-Precipitation bin, Figure 3a,d) for both shipping lanes.
Since the IMO regulation (Post-IMO), both shipping lanes exhibit significantly weaker lightning enhancements across most CAPE-Precipitation regimes (Figure 3b,e). The bin-by-bin differences between the Pre and Post-IMO lightning enhancement histograms for each shipping lane are shown in Figure 3 (c and f). For the vast majority of CAPE-Precip bins with statistically significant differences, the lightning enhancement Post-IMO has been weaker than that Pre-IMO. On average across all CAPE-Precip conditions, the lightning enhancement has decreased by 76% and by 47% for the Indian Ocean and South China Sea shipping lanes, respectively.(Figure 3c,f). Thus, for any given thunderstorm sampled, there is less lightning over the shipping lane since 2020 relative to the same type of thunderstorm—i.e., having the same associated CAPE and Precipitation—outside of the shipping lane.
@@
The decline in the lightning enhancement since 2020, which we have illustrated is present even after controlling for potential natural changes in convective activity or intensity and for relatively constant ship traffic, is therefore most consistent with a change in emissions composition in 2020 associated with the IMO regulation. If decreasing sulfur emissions over the shipping lanes has reduced the total number of viable CCN and disrupted an associated mechanism for lightning enhancement, then there should be a corresponding change in warm cloud microphysics. To further test this hypothesis, we use Moderate Resolution Imaging Spectroradiometer (MODIS) satellite observations of cloud droplet number (Nd) in warm (shallow) clouds over the Indian Ocean shipping lane, where the influence of land is weaker and ship emissions are stronger, during the high-lightning season. The retrievals of Nd follow the method outlined in Zhu et al., (2018), such that we sample only the brightest, most active warm clouds to ensure that assumptions of adiabatic ascent are as realistic as possible (see Appendix A). While we have also examined satellite observations of aerosol optical depth (AOD) over the region, AOD provides limited insights into CCN concentration changes in this region, and especially over the shipping lanes, due to the small size of ship-emitted particles and consequent marginal contributions to AOD (see SI).
Due to the optical thickness of DCC, the retrievals of Nd can only be done for shallow warm clouds. As a result, this analysis samples a different set of conditions than the lightning observations which arise essentially only from DCC. We thus assume that the behavior of Nd in shallow cumulus clouds from the same region is related to, though not necessarily a direct proxy for, Nd at cloud base in DCC. Unlike WWLLN lightning observations, which are influenced by the frequency of deep convection and therefore generally trend toward the ITCZ, the Nd retrievals mostly represent perturbations to CCN below shallow cumulus.
@@
In Figure 4, we show that prior to the IMO regulation, there is a clear trend in Nd toward land (north), as aerosol number concentrations are typically greater over land than over oceans (Figure S6), as well as a clear perturbation in Nd over the shipping lane. This Nd perturbation is roughly 10-15% above the average of droplet concentrations 150km north and 150km south, which is similar in magnitude to the shipping lane perturbations to Nd detected by Diamond et al., (2020) in Southeast Atlantic stratoculumus clouds. That the Nd perturbation over the Indian Ocean shipping lane is similar to that in other locations is a significant finding on its own, as observations of persistent, mean-state Nd perturbations by ships are rare (Diamond and Wood,, 2020). Since the IMO regulation, the Nd away from the shipping lane mostly maintain their previous levels, as indicated by the overlap in the 95% confidence intervals (shading), particularly to the north. We have partially accounted for variability in aerosol particles using MERRA-2 reanalysis estimates of continental dust and biomass burning aerosol variations across the region in constructing Figure 4. That said, a dearth of aerosol observations in the tropical Pacific likely limit the ability of reanalysis products to capture the full variability in CCN sources, possibly explaining some remaining differences in Nd pre and post IMO over the southern region of the domain.
@@
Most notably in Figure 4, the enhancement in Nd over the shipping lane has become essentially undetectable. The decline in Nd over the shipping lane establishes additional support for the declining lightning enhancement being related to a shift in aerosol particle number-size distributions, and consequently changes to the CCN distributions over the shipping lanes induced by the IMO regulation. It is important to note that Nd derived from shallow cumulus clouds will not be directly proportional to the CCN available for activation in deep convective clouds (DCC), nor to the lightning enhancements that might result from CCN enhancements. Given the strong updrafts that occur in DCC, aerosol particles can experience very high super-saturation with respect to liquid water, enabling smaller, often more numerous ultrafine aerosol (Hobbs et al.,, 2000) to act as CCN in DCC than can do so in shallow cumulus . The change in Nd is an otherwise rare, observable indication that the IMO regulations have clearly shifted aerosol particle distributions over the shipping lanes of interest here.
@@
We find that a previously identified enhancement in lightning stroke density over the two major shipping routes near the Port of Singapore has declined by over 40% since 2020, when the IMO regulation aimed at reducing sulfur emissions from maritime shipping came into effect. The decline of lightning is evident after controlling for natural variations in environmental conditions that characterize the intensity and frequency of thunderstorms, presented most clearly in Figure 3. While the ships themselves do likely act as attractors of lightning due to their prominence over the flat ocean (Peterson,, 2023), there is not evidence of a significant change in the number of ships traversing these shipping lanes over this time period (Figure S1). Further, that there was an independently observed perturbation to Nd over the Indian Ocean shipping lane prior to the IMO regulation that has since nearly vanished is indicative of a coincident change in CCN over the region. The statistically robust decline of the lightning enhancement, alongside the onset of the IMO regulation and observed changes in warm cloud microphysics, suggest that the mechanism for the lightning enhancement is likely induced by ship emissions of aerosol particles and precursors. The concomitant decline of Nd and lightning shown here lends credence to the set of CCN-mediated hypotheses previously proposed for invigoration of lightning (Fan et al.,, 2018; Abbott and Cronin,, 2021; Mansell and Ziegler,, 2013; Grabowski and Morrison,, 2020).
Precisely how increases in Nd might lead to enhanced lightning remains unresolved. Both the pre-IMO and post-IMO perturbations to Nd are smaller than the enhancements of lightning. In part, this difference may be explained by the different supersaturation conditions sampled by the Nd and lightning retrievals (warm, trade cumulus and deep convective environments, respectively), which will lead to different portions of the aerosol size distribution that can be activated in each case. Additionally, there may be non-linear relationships between Nd, secondary ice processes, and resulting charge separation; or additional interactions between DCC and aerosol, such as ice nucleation. Thornton et al. showed that DCC over the shipping lane exhibited anomalously higher radar reflectivity in the mixed phase region prior to 2015. While we find a similar indication for the four years prior to the IMO (see Figure S5) using a different satellite-based radar sensor, statistically robust tests of whether there have been changes in DCC reflectivity since the IMO regulation are not yet possible. However, we expect the statistical strength for the comparison between the two periods to increase, enabling a broader assessment of the post-IMO DCC response to maritime shipping emissions in the coming years.
@@
In situ observations of aerosol size distributions, needed to assess how the IMO regulations have led to changes in CCN in these regions, are notably lacking. Seppälä et al., (2021) find that, in the case of a factor of 10 reduction in ship fuel sulfur content, in-plume aerosol size distributions generally shift toward lower sizes and lower total number concentrations (i.e., that the number of ultrafine particles smaller than 50nm increases and the number of particles larger than 50nm decreases substantially). According to the ultrafine invigoration hypothesis proposed inFan et al., (2018), the increase in small ultrafine particles observed by Seppälä et al., (2021) would buffer enhancement losses of larger CCN, which is inconsistent with the strong decrease in lightning stroke density that we observe. Therefore, we conclude that the lightning enhancement over the shipping lane prior to the IMO regulation was mostly the result of enhancements in relatively larger aerosol particles (e.g. >50 nm in size) that perturbed: 1) cloud microphysics, such as elevated supercooled liquid water concentrations or rime splintering (Mansell and Ziegler,, 2013) and/or 2) updraft frequency, by means of heightened free tropospheric humidity or a mesoscale circulation response (Blossey et al.,, 2018; Grabowski and Morrison,, 2020). Enhancements in smaller ultrafine particles may play a role in the small, remaining lightning enhancement that has persisted over the shipping lanes post-IMO.
@@
We use the IMO regulation of ship fuel sulfur to identify and assess connections between maritime shipping emissions associated with international trade, pollution aerosol size and composition, DCC microphysics, and lightning stroke density. Further work combining in situ and remote sensing of aerosol and cloud microphysics together with lightning frequency is needed to clarify the mechanisms behind these connections, and to quantify the relative roles of dynamic and microphysical responses. Our findings herein show that these regions remain a useful testbed for understanding aerosol pollution impacts on deep convective clouds and lightning.

### v4

@@ Introduction
By acting as cloud condensation nuclei (CCN), aerosol particles influence clouds and, in turn, the Earth’s energy balance. Aerosol-cloud interactions represent a significant uncertainty in our understanding of the Earth’s climate (Boucher et al.,, 2013). Maritime ship traffic leads to the emission of aerosol particles and associated precursors into relatively clean marine air. These emissions enable study of how increased CCN perturb low-level marine stratus cloud droplet number distributions and related cloud macrophysical properties, such as cloud albedo and lifetime.(Diamond et al.,, 2020; Yuan et al.,, 2022; Durkee et al.,, 2000; Radke et al.,, 1989).
Deep convective cloud (DCC) systems occur throughout the tropics, and are essential to the Earth’s water and energy cycles (Feng et al.,, 2021). However, there is little consensus on the mechanisms or magnitudes of aerosol particle impacts on DCCs (Tao et al.,, 2012; Seinfeld et al.,, 2016; Igel and van den Heever,, 2021; Varble et al.,, 2023). Thornton et al., (2017) documented a potential case of persistent maritime aerosol-DCC interactions analogous to stratocumulus ship tracks, with the discovery of enhancements in lightning over major shipping lanes passing through the Indian Ocean and South China Sea (Fig. 1).
@@
Several mechanisms have been proposed to explain how aerosol particles from ship emissions could enhance lightning frequency, all of which involve enhanced cloud droplet nucleation (Twomey,, 1977), leading to either 1) perturbations to super-cooled liquid water and ice hydrometeor distributions and enhanced charge separation in the mixed-phase region of DCC (Mansell and Ziegler,, 2013; Blossey et al.,, 2018; Sun et al.,, 2024; Takahashi and Miyawaki,, 2002; Deierling et al.,, 2008); or 2) an increase in the frequency or intensity of deep convection due to changes in the vertical distribution of humidity (Abbott and Cronin,, 2021) or heating (Fan et al.,, 2018; Grabowski and Morrison,, 2020). Some combination of 1 or 2 is also possible.
@@
In January 2020, the International Maritime Organization (IMO) reduced the amount of allowable sulfur in fuel by a factor of seven, from 3.5% to 0.5% to curb effects of maritime shipping on air pollution (IMO,, 2020). Recent analyses of shallow stratocumulus marine clouds over shipping lanes find changes to cloud brightness, droplet number, and droplet size associated with the IMO regulation, presumably due to the shift in aerosol size distribution (Watson-Parris et al.,, 2022; Yuan et al.,, 2022; Diamond,, 2023).
@@
We investigate whether the IMO fuel sulfur regulation has impacted lightning over the shipping lanes in the tropical Indian Ocean and South China Sea. We find the shipping lane lightning enhancement decreases significantly with the onset of the IMO regulation and that this decrease persists across a range of atmospheric conditions. We further show that the mean cloud droplet number concentration of shallow warm clouds over the Indian Ocean shipping lane was enhanced before the IMO regulation and also exhibits a decrease since the IMO regulation. We discuss the implications of these new results for mechanisms of shipping lane lightning enhancement and aerosol-DCC interactions.
@@
The Port of Singapore accounts for  20% of the world’s bunkering fuel demand. The two primary shipping lanes it services—the Indian Ocean and South China Sea (hereafter ”the shipping lanes”)—have nearly an order of magnitude higher traffic than other shipping lanes around the world (Figure 1, top panel) (Mao et al.,, 2022). As shown in Figure 1 (middle panel), prior to 2020, the mean absolute lightning stroke density measured by the World Wide Lightning Detection Network (WWLLN) remains enhanced over these shipping lanes, consistent with Thornton et al 2017. Since 2020, however, when the IMO regulation of sulfur emissions began, absolute lightning stroke density over the shipping lanes has declined by about 1 stroke km-2 year-1 as indicated by the difference between the means of years 2020-2023 and 2010-2019 (Figure 1, bottom panel). While some of the largest absolute declines in lightning since 2020 occur over the shipping lanes, lightning has increased or decreased in other parts of this region as well. As we illustrate below, variability in the dynamic and thermodynamic context for convection over these shipping lanes, as well as potential changes to shipping due to pandemic and other disruptions to trade must be taken into account to better isolate the potential impacts of shipping emissions.
Ship traffic in the vicinity, as measured by vessel fuel sales at the Port of Singapore, has been relatively constant or even increased since 2020 (Figure S1) (Port of Singapore,, 2024). The disruption by COVID-19 did not obviously decrease activity at the port, seeming only to have briefly slowed the growth of cargo throughput for 2-3 months in 2020 (Gu et al.,, 2023). Therefore, we focus on controlling for the variability in background meteorological conditions that impact the frequency and intensity of convection, and thus lightning, over the shipping lanes.
@@
We first examine the shipping lane lightning enhancement using two controls on background meteorology: 1) we only sample data from times and pixels with precipitating clouds(Huffman et al.,, 2015; Pradhan and Markonis,, 2023; Watters et al.,, 2023); and 2) we restrict analyses to the specific seasons in each region favorable for lighting (November to April in the Indian Ocean; June to November in the South China Sea). Using these criteria, we composite lightning observations as a function of distance to the shipping lanes, the center of which we define as the peak in shipping emissions from the EDGAR emissions inventory (see Methods). As shown in Figure 2, mean absolute lightning exhibits a clear enhancement over the shipping lane before 2020 (pre-IMO), between approximately 150km south to 150km north of the shipping lanes, and that has decreased since the regulation onset in 2020 (post-IMO) (Figure 2a). Consistent with Figure 1, lightning in this context has fallen region-wide, particularly to the south of the shipping lanes. We next examine how variability in convective activity between the two periods (pre and post IMO) has affected the regional lightning changes.
@@
To account for inter-annual variability in the frequency and intensity of convection in the region, we regress the observed annual lightning at a given distance from the shipping lane against three variables known to relate to lightning frequency: Convective Available Potential Energy (CAPE, discussed further below) and precipitation rates (Romps et al.,, 2018), as well as the annual mean Oceanic Niño Index (ONI). The influence of El Niño is assumed to be uniform across the domain for a given year. Inter-annual variability in the MJO was small and had a negligible impact when included in the regression (see SI). Further details on the regression are in Appendix A. The regressed variables explain  33% of variance of the annual means. We subtract the regressed lightning from the observed annual mean, leaving the anomalous mean lightning stroke density that cannot be explained by interannual variability in storm occurrence and intensity, shown in Figure 2b.
@@
The annual anomalous enhancement in lightning over the shipping lanes prior to 2020 is even more clear after regressing out meteorological variability, and so is the near step-change decrease in the anomaly after 2020 (2b). Prior to the IMO regulation, essentially 100% of fuel sold at the port was high-sulfur (2b, right axis); correspondingly, the lightning anomaly over the shipping lane was 4.0 strokes km-2 year-1 on average and was never below 3 strokes km-2 year-1 for more than one year at a time. Adoption of the IMO regulation was prompt in 2020, as indicated by the change in high-sulfur fuel from 100% to less than 35% of fuel sold at the Port of Singapore. The Port of Singapore experienced little attenuation of fuel sales at the onset of COVID-19 (Gu et al.,, 2023), and total fuel sales have increased since 2020 consistent with higher traffic (Figure S1). Since 2020, the shipping lane lightning enhancements compared to adjacent regions have declined by  50% to  2 strokes km-2 year-1 on average.
@@
To further control for higher-frequency variations in convective activity and intensity, we examine the lightning enhancement in a 2-dimensional CAPE and precipitation space, using 3-hourly coincident observations of CAPE, precipitation, and lightning. Romps et al., (2018) showed that CAPE $\times$ precipitation is a reasonable proxy for lightning frequency over land, and Cheng et al., (2021) showed that with a CAPE threshold, it is a reasonable proxy for tropical oceanic lightning frequency. The 3-hourly CAPE and precipitation observations implicitly capture variability arising from more indirect sources, such as sea surface temperatures (SST), MJO, fronts, etc (see SI for further discussion).
We compute lightning frequency in each CAPE-Precipitation bin using data from a region centered over each shipping lane and from reference regions adjacent to the shipping lanes, similar to those in (Thornton et al.,, 2017) (see Figure S2). We then compute a relative enhancement in lightning over the shipping lanes, before and after the onset of the IMO regulation, by taking the difference between corresponding CAPE-Precipitation bin-means in the shipping lane and associated reference box. The resulting shipping lane lightning enhancements as a function of both CAPE and Precipitation are shown in Figure 3. Before the IMO regulation (Pre-IMO), a shipping lane lightning enhancement existed in nearly every thermodynamic setting (e.g., in each CAPE-Precipitation bin, Figure 3a,d) for both shipping lanes.
Since the IMO regulation (Post-IMO), both shipping lanes exhibit significantly weaker lightning enhancements across most CAPE-Precipitation regimes (Figure 3b,e). The bin-by-bin differences between the Pre and Post-IMO lightning enhancement histograms are shown in Figure 3 (c and f). On average across all CAPE-Precip conditions, the lightning enhancement has decreased by 76% and by 47% for the Indian Ocean and South China Sea shipping lanes, respectively (Figure 3c,f). Thus, for the same convective setting characterized by CAPE and Precipitation rates, lightning stroke density over the shipping lanes is significantly less enhanced compared to adjacent regions after 2020 than it was prior to 2020.
@@
Based upon the above, we hypothesize that the decline in the lightning enhancement since 2020 is most consistent with the IMO regulation changing CCN in the region. If decreasing sulfur emissions over the shipping lanes has reduced the total number of viable CCN and disrupted an associated mechanism for lightning enhancement, then there should be a corresponding change in warm cloud microphysics. To further test our hypothesis, we use Moderate Resolution Imaging Spectroradiometer (MODIS) satellite observations of cloud droplet number (Nd) in warm (shallow) clouds over the Indian Ocean shipping lane, where the influence of land is weaker and ship emissions are stronger, during the high-lightning season. The retrievals of Nd follow the method outlined in Zhu et al., (2018), such that we sample only the brightest, most active warm clouds to ensure that assumptions of adiabatic ascent are as realistic as possible (see Appendix A).
Due to the optical thickness of DCC, the retrievals of Nd can only be done for shallow warm clouds. As a result, this analysis samples a different set of conditions than the lightning observations which arise essentially only from DCC. We thus assume that the behavior of Nd in shallow cumulus clouds from the same region is related to, though not necessarily a direct proxy for, Nd at cloud base in DCC. We have partially accounted for variability in aerosol particles from continental dust and biomass burning aerosol using MERRA-2 reanalysis estimates across the region in constructing Figure 4 to better isolate the effect of shipping emission changes on Nd. Limited observations in the region likely hinder the ability of reanalysis products to capture the full variability in CCN sources (see SI for discussion of aerosol optical depth), possibly explaining some remaining differences in Nd over the southern region of the domain pre and post 2020.
@@
In Figure 4, we show that prior to the IMO regulation, there is a clear trend in Nd toward land (north), as aerosol number concentrations are typically greater over land than over oceans (Figure S6), as well as a clear perturbation in Nd over the shipping lane. This Nd perturbation is roughly 10-15% above the average of droplet concentrations 150km north and 150km south, which is similar in magnitude to the shipping lane perturbations to Nd detected by Diamond et al., (2020) in Southeast Atlantic stratoculumus clouds. The Nd perturbation over the Indian Ocean shipping lane is a significant finding on its own, as observations of persistent, mean-state Nd perturbations by ships are rare (Diamond and Wood,, 2020), especially for convectively active regions we show here.
@@
Since the IMO regulation, the Nd away from the shipping lane mostly maintain their previous levels, as indicated by the overlap in the 95% confidence intervals (shading), particularly to the north. Meanwhile, the enhancement in Nd over the shipping lane has become essentially undetectable 4. The decline in Nd over the shipping lane relative to the surrounding region establishes additional support for the declining lightning enhancement being related to a shift in aerosol particle number-size distributions, and consequently changes to the CCN distributions over the shipping lanes induced by the IMO regulation. Nd derived from shallow cumulus clouds will not be directly proportional to the CCN available for activation in deep convective clouds (DCC) due to the different supersaturation conditions present in each cloud type ((Hobbs et al.,, 2000), nor to the lightning enhancements that might result from CCN enhancements. However, the change in Nd is an additional observable indication, independent of the lightning observations, that the IMO regulations have clearly shifted aerosol particle distributions over the shipping lanes of interest here.
@@
We find that a previously identified enhancement in lightning stroke density over the two major shipping routes near the Port of Singapore has declined by over 40% since 2020, when the IMO regulation of maritime shipping sulfur emissions came into effect. The decline is evident after controlling for natural variations in environmental conditions that characterize the convection intensity and frequency (see Figure 3). While ships may act as lightning attractors (Peterson,, 2023), there is not evidence of a change in the number of ships traversing these shipping lanes over this time period (Figure S1). Further, an independently observed perturbation to Nd over the Indian Ocean shipping lane prior to the IMO regulation has since nearly vanished, indicating a coincident change in CCN over the region. The concomitant decline of Nd and lightning, timed with the onset of compliance with the IMO regulation, provides new evidence to support the set of CCN-mediated hypotheses previously proposed for invigoration of lightning (Fan et al.,, 2018; Abbott and Cronin,, 2021; Mansell and Ziegler,, 2013; Grabowski and Morrison,, 2020).
Precisely how increases in CCN and Nd caused by ship emissions might lead to enhanced lightning remains unresolved. Both the pre-IMO and post-IMO perturbations to Nd are smaller than the enhancements of lightning, which may be explained by high scavenging rates in DCC during the high lightning season, as well as the different supersaturation conditions sampled by the Nd (shallow cumulus) and lightning retrievals (deep convection), non-linear relationships between Nd and ice secondary ice production, and ice nucleation.
@@
Seppälä et al., (2021) find that a factor of 10 reduction in ship fuel sulfur content shifts emitted aerosol particles to smaller sizes and lower total number concentrations. Ultrafine particles <50nm increase and those >50nm decrease substantially, which implies that ultrafine particle invigoration of updraft velocities proposed in Fan et al., (2018) was not contributing to the shipping lane lightning enhancement pre–IMO. We conclude the lightning enhancement prior to the IMO regulation was mostly the result of higher concentrations of larger aerosol particles (e.g. >50 nm) that perturbed: 1) cloud microphysics, such as elevated supercooled liquid water concentrations or rime splintering (see SI) (Mansell and Ziegler,, 2013); and/or 2) updraft frequency, by means of heightened free tropospheric humidity or a mesoscale circulation response (Blossey et al.,, 2018; Grabowski and Morrison,, 2020). Enhancements in ultrafine particles may play a role in the smaller non-zero shipping lane lightning enhancement that has persisted post-IMO.
@@
The IMO regulation of ship fuel sulfur illustrates connections between international trade, air pollution, DCC microphysics, and lightning. Further work combining in situ and remote sensing of aerosol and cloud microphysics together with lightning frequency is needed to clarify the mechanisms behind these connections, and to quantify the relative roles of dynamic and microphysical responses. Our findings herein show that these regions remain a useful testbed for understanding aerosol pollution impacts on deep convective clouds and lightning.